a study conducted by the pew research center reported that 58% of cell phone owners used their phones for purchase guidance. A random sample of 15 cell phone owners is studied to understand how cell phones are used to make purchase decisions. define the random variable and find the mean and standard deviation
Answer #1

Solution:

Random variable is a variable whose possible values outcomes are numerical outcomes of a random phenomenon.

This is an example of discrete random variable as we can count these number rather than measure.

Mean of distribution = no. Of sample * probability of success = 15*0.58 = 8.7

Standard deviation of distribution = sqrt(no.of sample*probability of success *probability of failure) = sqrt(15*0.58*0.42) = sqrt(3.654)= 1.91